# To Make, or to Take, That Is the Question: Impact of LOB Mechanics on Natural Trading Strategies

### v1

@@ Abstract
Working at a very granular level, using data from a live trading experiment on the Binance linear Bitcoin perpetual—the most liquid crypto market worldwide—we examine the effects of (i) basic order book mechanics and (ii) the strong persistence of price changes from the immediate to the short timescale, revealing the interplay between returns, queue sizes, and orders’ queue positions. For maker orders, we find a negative correlation between fill likelihood and subsequent short-term returns, posing a significant challenge for maker order-based strategies, while the main hurdle with taker orders is overcoming the taker fee. These dynamics render natural (and commonly-cited) trading strategies highly unprofitable. Finally, we use the understanding gained to identify situations (Reversals) in which a successful trading strategy can operate; we construct a signal for Reversals and demonstrate its efficacy.
Keywords market making, limit order book, fill probability, optimal execution, adverse selection, high-frequency trading
@@ 1 Introduction
Market participants in LOBs can execute trades as (liquidity) takers, makers (liquidity providers), or as a combination of both. Maker orders (that is, passive limit orders posted at the top bid/ask price, or deeper into the book) [^1] In then academic literature, ‘limit order’ often refers exclusively to same-side maker orders (passive buy orders on the bid side, sell orders on the ask side). In practice, limit orders can also be posted on the ‘other’ side of the book, as marketable limit orders, for example a buy limit order posted at the top ask price. Such an order executes immediately as a taker order (paying the taker fee) against any passive orders from the top of the book to the limit price, with the remainder forming a new queue of one maker order at the limit price. We adopt the more precise terminology where a limit order can act as a maker or taker order, or partially fill as a taker with the remainder left open as a maker order, and we encourage other academics to do the same. offer a better price (by at least the spread) and significantly lower trading fees, sometimes even earning a rebate. These advantages may suggest that maker orders are a more cost-effective way to execute trades than taker orders (market orders or marketable limit orders). However, they also come with two notable disadvantages.
@@
First, maker orders face execution uncertainty: they do not fill immediately and may not fill at all, resulting in a fill probability less than 1; market orders have a fill probability of 1.222Marketable limit orders have a fill probability slightly below 1, see  Cartea et al. [2021]. Second, there is a negative correlation between a maker order’s fill probability and its subsequent return conditioned on filling. This effect, termed the negative drift of maker orders by DeLise [2024], is especially pronounced at the short timescales which are the primary focus of our paper. A simple mechanical fact underpins this observation. If the next price move is against a maker order in the top-of-book queue, that order automatically fills, with probability 1, a phenomenon sometimes referred to as adverse selection. Such price moves are often driven by short-lived arbitrage opportunities, where prices adjust with a lag to new information, exposing makers to the risk of trading at stale quotes (termed toxic arbitrages in Foucault et al. [2016b]). If, in contrast, the next price change is both in the same direction as that of the maker order and occurs before the order is filled the order may no longer be at the top of the book, thus not profiting from the favorable price change. In summary, it is not easy, when using a maker order, to achieve both a high fill probability and a positive subsequent return if filled. This connection between an order’s terminal state (fill-or-not) and the price trajectory is underappreciated and largely ignored in the academic literature.
@@
Market orders, in contrast, fill with probability 1 (although a large order may still walk the book, and thereby cause an immediate price change as above). Moreover, simple strategies such as trading in the direction of the order imbalance between the top bid and ask queues (formally defined in (1) below) have been shown to yield positive short-term returns Lipton et al. [2013], Stoikov [2017]. The challenge to the liquidity-taking trader is to achieve positive returns exceeding the exchange’s taker fee, which can be viewed as the profitability threshold.
@@
We address the questions of fill probability and price drift of maker orders posted at the best prices (top bid and top ask); these price levels are of paramount importance for practitioners because the majority of trading volume (both maker and taker) occurs at them. We also note that, on Binance, the spread is virtually always one tick wide, except fleetingly after price changes, whereupon there is fierce competition to close the spread immediately. Although our analysis is specific to the Bitcoin market, our approach is broadly applicable and our findings should have wider relevance.
@@ 1.1 Related Literature and Context
In our search for answers to the above questions, our paper traverses a lot of territory; we begin with a high-level overview of relevant topics and related literature, before proceeding to a more detailed analysis in the sections that follow.
@@ 1.1.1 Maker Order Execution
The study of fill probabilities for maker orders has garnered significant interest in the literature. One should distinguish between estimating (experimentally) the fill probability of randomly placed orders, agnostic to any trading or cancellation signal, and inferring (from historical LOB data) the fill probability of actual orders placed by market participants. The former is a more fundamental question: if one were to post an order, perhaps conditioned on the order book state or other observable information, how likely is that order to fill? The latter question examines the behavior of market participants, determining the fill frequency of the orders they chose to place. While both are of interest, the practical applicability of the latter is diminished by systematic biases, which we will refer to as signal bias: the estimated fill probabilities are implicitly conditioned on some trader’s decision to place and not cancel the order, both of which may depend on private, unobservable signals that determine the trader’s actions. Two recent works, Fabre and Ragel [2023] and Lokin and Yu [2024], address the latter question concerning the behavior of market participants, albeit without explicitly highlighting the important distinction explained above. Fabre and Ragel [2023] conclude that fill probabilities are significantly influenced by market conditions such as bid-ask spread, order size, and liquidity imbalance. Meanwhile, Lokin and Yu [2024] develop a stochastic model of LOB dynamics as a state-dependent queueing system and estimate fill probabilities of maker orders within that framework, concluding that this can be done to a reasonable degree of accuracy. The estimated fill probabilities can differ significantly depending on which of the two questions one seeks to answer; for instance, Lokin and Yu [2024] reports a fill probability of less than 2% for orders placed by other market participants when the spread is one tick wide and bid and ask queues are balanced, compared to the 50% fill probability we find for orders placed agnostic to any trading signal in the same situation.
@@
The primary motivation outlined in the strand of literature addressing the more fundamental question is the trade-off between maker and taker orders: maker orders offer better prices and fees [^4] Though, as far as we are aware, the fee improvement is left unmentioned in the existing literature. but do not execute immediately and may fail to execute at all. A number of recent papers have explored this topic. In Maglaras et al. [2021], the authors simulate synthetic top-of-book maker orders on the NASDAQ exchange and record whether or not they fill, as well as their time-to-fill, given actual market activity after the submission of the synthetic maker order. They then construct a deep learning-based model to predict fill probabilities and the conditional expected fill times, which outperforms benchmarks. The more recent work by Arroyo et al. [2023] is similar in spirit but differs in methodology: they employ convolutional neural networks and transformers to perform survival analysis of maker orders, including ones posted at deeper levels in the order book. However, they do not distinguish the side of the book in which the order is posted, leading to inaccurate estimates when comparing the fill probabilities of two opposite-side orders posted at the same depth in a skewed order book imbalance (one which is bid-dominated or ask-dominated).
@@
We add to this strand of literature as follows. Our data set of actual orders and their outcomes is free from signal-bias, and unlike a synthetic-order approach, not reliant on a-posteriori inference of when such a synthetic order might have filled. [^5] Such inferences may be possible with a reasonable degree of accuracy, but they are still open to error, in some cases due to factors such as latency. We leverage our data to analyze fill probabilities of maker orders, and demonstrate that they can be described in simple terms, and indeed be predicted with a high degree of accuracy, as a function of top-of-book quantities and an order’s queue position, calling into question the need for black-box machinery lacking interpretability and ease of replicability to solve what turns out to be a relatively simple problem.
@@ 1.1.2 Trade-offs and Trading Decisions
As well as using our data set to estimate fill probabilities, we use it to illustrate the difficulty of leveraging that predictability into trading edge. In particular, in addition to the trade-off between price and immediacy of execution, we highlight the trade-off between fill probability itself and the subsequent returns of filled orders. We support our earlier assertions: first, that it is not at all easy to combine a high fill probability for maker orders with a positive subsequent return, due to one or other of adverse selection if filled, and low fill probability when subsequent returns are good; and second, for taker orders, to achieve short-term returns in excess of the taker fee. One can see this as an instance of an ‘Unprofitability Principle’ stated in cartoon form as
Ease of prediction×ease of $\text{Ease of prediction}\times\text{ease of exploitation}<c$
for some positive constant $c$. As far as we are aware, only a few papers in the literature provide detailed analyses of adverse selection and outcomes of maker orders, including the recent work  DeLise [2024], with which our findings and conclusions align. Earlier contributions such as  Foucault et al. [2003] discuss the need for market makers to monitor information sources to cancel stale quotes, while  Foucault et al. [2016a] and  Foucault et al. [2016b] examine the adverse selection risks market makers face from informed HFT takers reacting to news or arbitrageurs exploiting short-lived price adjustments; the data gathered through our live trading experiments builds on and supports these earlier findings.
@@ 1.1.3 Trading with Signals
We have noted (and analyze in more detail below), that some apparently natural strategies drawing on basic publicly available information (for example, posting a taker order in the direction of the imbalance or a maker order with high fill probability) may not yield a trading advantage. A trader may, however, take a more positive view if they have extra information in the form of a private signal, for example algorithmically generated.
Suppose, for example, that the top bid queue is large and the top ask queue is small—a situation we describe as a price-positive imbalance. Intuitively, it is then likely that the shorter (ask) queue is the first to empty, in which case the spread widens and may be refilled with a new queue on the bid side, giving a price increase. This argument is corroborated in practice: it is well known that imbalance is positively correlated with the direction the next price movement. (For further insights, refer to  Gould and Bonart [2015].) Now suppose that a trader supplements this modestly return-positive signal with a convincing private signal, suggesting an enhanced short-term return. How best to capitalize on this opportunity? Submitting a maker order at the end of the large bid queue is unlikely to result in a fill, thus not profiting from the prediction (in the less frequent instances where it does fill, the short-term return is likely to be negative, as we discuss in more detail in Section 5). Conversely, a taker order is certain to fill and benefit from any favorable price change; but they must pay the taker fee.
@@
These arguments illustrate the important point that the choice of order type and execution mode depends on the kind of signal in response to which the trader makes decisions. The literature contains numerous discussions of trading with signals, often under the umbrella term ‘optimal market making’ when they involve maker orders. For instance, Cartea and Sánchez-Betancourt [2023] conclude that predictive signals, referred to as alpha signals, can boost market maker profitability. Their conclusions are arrived at under the assumption that maker orders fill upon the arrival of the next opposite-direction taker order; that is, the maker order is always at the front of the queue. As just argued, and discussed in more detail later, this is not generally tenable. Other papers such as  Fushimi and Rojas [2018] and  Ait-Sahalia and Saglam [2013] rely on the same assumption, while others still make different assumptions which are also problematic: for example, Spooner et al. [2018] assumes that cancellations remove liquidity uniformly from the queue.
@@
Our examples above further illustrate the influence of queue position—whether submitting at the end of a long queue or a short one—on an order’s fill probability and the effectiveness of using maker orders to act on a trading signal. For example, posting maker orders in the direction of the imbalance implies posting at the back of the longer of the bid- and ask queues, unless one is the first to close the spread after a price change, which is extremely competitive and typically involves latency optimization and/or special order types. Conversely, one can obtain a good queue position (in the shorter of the two queues) by posting a maker order opposite to the imbalance-implied movement, but this entails trading against the expected return. Although queue position is rather sparsely addressed in the academic literature, Moallemi and Yuan [2016] find that bad queue positions are associated with lower fill probabilities and increased adverse selection, findings that our results corroborate; additionally, Donnelly and Gan [2018] explore the interplay of queue position, adverse selection, and cancellation behavior, showing that within their modeling framework, the optimal strategy sometimes involves submitting an order counter to the imbalance to benefit from a queue position close to the front-of-queue, a result that is in agreement with our findings.
@@ 1.2 Notation and Definitions
Our analysis is set in a standard LOB, a data structure that records queues of maker orders waiting to be executed or canceled, prioritized by price and then by time of arrival, and quantized by the minimum order size, at price levels incremented by the tick size, The state of the order book at time $t$ is defined by the bid and ask sides:
@@
Its value is close to $+1$ when the liquidity is significantly greater at the top bid level than at the top ask level, and conversely for values near $-1$. As noted above, these extremes are associated with price increases and decreases respectively.
We focus exclusively on those maker orders posted at the top level (and taker orders filling them), meaning sell orders at the top ask price $p^{a}_{t,1}$ and buy orders at the top bid price $p^{b}_{t,1}$. These two order book levels hold a special significance as a large fraction of overall trading volume is executed against orders posted at the touch, making them especially important for market makers. For brevity, we shall use the term maker order to mean exclusively a maker order posted posted at one of the two top levels. The analysis of maker orders posted at deeper levels in the LOB requires different tools and is beyond the scope of this paper.
@@
Consider a buy order (and analogously for sell orders) $\mathfrak{o}$ posted at the top bid price $p^{b}_{t,1}$ at time $t$, indexed by $j_{0}$ in the sequence $\left(q^{b}_{t,1,j}\right)_{j}$ of individual order quantities. As illustrated in Figure 1, we define the liquidity ahead of the order as the total quantity if all preceding orders in the queue:
@@
and equal to 00 when there are no other orders in the queue ($\operatorname{LA}_{t}=\operatorname{LB}_{t}=0$). An order whose queue position is 00 is at the front of the queue; if it is $1$, the order is at the back.
@@
Consider an order $\mathfrak{o}$ (either maker or taker) posted at time $t_{0}$ when the top-of-book price is $p_{t_{0}}$.  [^6] Conventionally, for a maker order, $p_{t_{0}}$ is the top price on the near-side, while for a taker order, it is the top price on the opposite-side. We define a sign function, $\operatorname{sgn}(\mathfrak{o})$, equal to $+1$ for buy orders and $-1$ for sell orders, and multiply the drift values by this sign to obtain a direction-independent measure, with positive (negative) values indicating favorable (adverse) drift.

### v2

# The Market Maker’s Dilemma: Navigating the Fill Probability vs. Post-Fill Returns Trade-Off

@@ Abstract
Using data from a live trading experiment on the Binance Bitcoin perpetual, we examine the effects of (i) basic order book mechanics and (ii) the persistence of price changes from immediate to short timescales, revealing the interplay between returns, queue sizes, and orders’ queue positions. We document a fundamental trade-off: a negative correlation between maker fill likelihood and post-fill returns. This dictates that viable maker strategies often require a contrarian approach, counter-trading the prevailing order book imbalance. These dynamics render commonly-cited strategies highly unprofitable, leading us to model ‘Reversals’: situations where a contrarian maker strategy at the touch proves effective.
The Market Maker’s Dilemma:
Navigating the Fill Probability vs. Post-Fill Returns Trade-Off
Jakob Albers1 Mihai Cucuringu1,2,3 Sam Howison4 Alexander Y. Shestopaloff5,6
1Department of Statistics, University of Oxford, Oxford, UK
2Oxford-Man Institute of Quantitative Finance, University of Oxford, UK
3Department of Mathematics, University of California Los Angeles, Los Angeles, US
4Mathematical Institute, University of Oxford, Oxford, UK
5School of Mathematical Sciences, Queen Mary University of London, London, UK
6Department of Mathematics and Statistics, Memorial University of Newfoundland, St. John’s, NL, Canada
Corresponding author: jakob.albers@merton.ox.ac.uk
November 23, 2025
Keywords: market making, limit order book, fill probability, optimal execution, adverse selection, high-frequency trading
JEL classification: G13, G14, C81, C90
@@ 1  Introduction
Market participants in LOBs can execute trades as (liquidity) takers, makers (liquidity providers), or as a combination of both. Maker orders (that is, passive limit orders posted at the top bid/ask price, or deeper into the book) [^1] In the academic literature, ‘limit order’ often refers exclusively to same-side maker orders (passive buy orders on the bid side, sell orders on the ask side). In practice, limit orders can also be posted on the ‘other’ side of the book, as marketable limit orders, for example a buy limit order posted at the top ask price. Such an order executes immediately as a taker order (paying the taker fee) against any passive orders from the top of the book to the limit price, with the remainder forming a new queue of one maker order at the limit price. We adopt this more precise terminology—where a limit order can act as a maker, a taker, or a combination—and encourage other academics to do the same. offer a better price (by at least the spread) and significantly lower trading fees, sometimes even earning a rebate. These advantages may suggest that maker orders are a more cost-effective way to execute trades than taker orders (market orders or marketable limit orders). However, they also come with two notable disadvantages.
@@
First, maker orders face execution uncertainty: they do not fill immediately and may not fill at all, resulting in a fill probability less than 1; market orders have a fill probability of 1.222Marketable limit orders have a fill probability slightly below 1, see  cartea2021latencyrisk. Second, there is a negative correlation between a maker order’s fill probability and its subsequent return conditioned on filling. This effect, termed the negative drift of maker orders by DeLise:2024aa, is especially pronounced at the short timescales which are the primary focus of our paper. A simple mechanical fact underpins this observation. If the next price move is against a maker order in the top-of-book queue, that order automatically fills, with probability 1, a phenomenon often referred to as adverse selection Menkveld2013. Such price moves are often driven by short-lived opportunities, perhaps even amounting to arbitrage, arising as prices adjust with a lag to external signals, exposing makers to the risk of trading at stale quotes (termed toxic arbitrages in foucault_2016). If, in contrast, the next price change is both in the same direction as that of the maker order and occurs before the order is filled, then the order may no longer be at the top of the book, thus not profiting from the favorable price change. In summary, the fundamental challenge for maker strategies is to achieve both a high fill probability and a positive return if filled: the trade-off between those two aspects renders profitable maker trading inherently difficult. Few academic studies examine it in detail or propose ways to walk that tightrope; we see below that this may necessitate submitting orders under apparently adverse conditions using an additional signal (without which the orders would be money-losing).
@@
Market orders, in contrast, fill with probability 1 (although a large order may still walk the book, and thereby cause an immediate price change as above). Moreover, simple strategies such as trading in the direction of the order imbalance between the top bid and ask queues (formally defined in (1) below) have been shown to yield positive short-term returns Lipton:2013aa; Stoikov_2017. The challenge to the liquidity-taking trader is to achieve positive returns exceeding the exchange’s taker fee, which can be viewed as the profitability threshold.
@@
We address the questions of fill probability and price drift of maker orders posted at the best prices (top bid and top ask); these price levels are of paramount importance for practitioners because the majority of trading volume (both maker and taker) occurs at them. We also note that, on Binance, the spread is virtually always one tick wide, except fleetingly after price changes, whereupon there is fierce competition to close the spread immediately.
@@ 1.1  Related Literature and Context
At first sight, we have posed relatively specific microstructural questions. However, being generic, they sit in, and contribute to, a far broader context than the Bitcoin market that we study. We therefore begin with a high-level overview of relevant topics and related literature, before proceeding to a more detailed analysis in the sections that follow.
@@ 1.1.1 Maker Order Execution
The study of fill probabilities for maker orders has garnered significant interest in the literature. One should distinguish between estimating (experimentally) the fill probability of randomly placed orders, agnostic to any trading or cancellation signal, and inferring (from historical LOB data) the fill probability of actual orders placed by market participants. The former is a more fundamental question: if one were to post an order, perhaps conditioned on the order book state or other observable information, how likely is that order to fill? The latter question examines the behavior of market participants, determining the fill frequency of the orders they chose to place. While both are of interest, the practical applicability of the latter is subject to systematic biases, which we will refer to as signal bias: the estimated fill probabilities are implicitly conditioned on the trader’s decision not to cancel the order, with the trader following a cancellation policy that is a priori unknown to us—and the importance of this can hardly be overstated given that >>95% of submitted maker orders are cancelled; see Gould:2010aa and Brogaard2014. Traders generally also follow private, a priori unknown order submission policies, e.g., submitting only when they deem adverse selection to be unlikely.
Two recent works, Fabre_2023 and Lokin:2024aa, use the latter approach concerning the behavior of market participants, albeit without explicitly highlighting the important distinction explained above. Fabre_2023 conclude that fill probabilities are significantly influenced by market conditions such as bid-ask spread, order size, and liquidity imbalance. Meanwhile, Lokin:2024aa develop a stochastic model of LOB dynamics as a state-dependent queueing system and estimate fill probabilities of maker orders within that framework, concluding that this can be done to a reasonable degree of accuracy. The estimated fill probabilities can differ significantly depending on which of the two aforementioned questions one seeks to answer; for instance, Lokin:2024aa reports a fill probability of less than 2% for orders placed by other market participants when the spread is one tick wide and bid and ask queues are balanced, compared to the 50% fill probability we find in the same situation. This is largely a cancellation effect: real makers cancel orders, biasing historical estimates low, while our no-cancellation policy yields a baseline unaffected by private signals (as discussed further in Section 6).
@@
The primary motivation outlined in the strand of literature addressing the more fundamental question is the trade-off between maker and taker orders: maker orders offer better prices and fees [^4] Though, as far as we are aware, the fee improvement is left unmentioned in the existing literature. but do not execute immediately and may fail to execute at all. A number of recent papers have explored this topic. In Maglaras_2021, the authors simulate synthetic top-of-book maker orders [^5] That is, fictitious orders ‘inserted’ at the back of the relevant queue, whose subsequent queue place and execution (or not) can be tracked using subsequent LOB-evolution data. Note, however, that, unlike real orders, they have no effect on the LOB dynamics. on the NASDAQ exchange and record whether or not they fill, as well as their time-to-fill, given actual market activity after the submission of the synthetic maker order. They then construct a deep learning-based model to predict fill probabilities and the conditional expected fill times, which outperforms benchmarks. The more recent work by Arroyo_2023 is similar in spirit but differs in methodology: they employ convolutional neural networks and transformers to perform survival analysis of maker orders, including ones posted at deeper levels in the order book. However, they do not distinguish the side of the book in which the order is posted relative to the imbalance, leading to inaccurate estimates when comparing the fill probabilities of two opposite-side orders posted at the same depth in a skewed order book imbalance (one which is bid-dominated or ask-dominated).
@@
We add to this strand of literature as follows. Our data set of actual orders and their outcomes is free from signal-bias, and unlike a synthetic-order approach, not reliant on a posteriori inference of when such a synthetic order might have filled. [^6] Such inferences may be possible with a reasonable degree of accuracy, but they are still open to error, in some cases due to factors such as latency. We leverage our data to analyze fill probabilities of maker orders, and demonstrate that they can be described in simple terms, and indeed be predicted with a high degree of accuracy, as a function of top-of-book quantities and an order’s queue position, calling into question the need for black-box machinery lacking interpretability and ease of replicability to solve what turns out to be a relatively simple problem.
@@ 1.1.2 Trade-offs and Trading Decisions
As well as using our data set to estimate fill probabilities, we use it to illustrate the difficulty of leveraging that knowledge (or predictability) into trading edge. In particular, in addition to the trade-off between price and immediacy of execution, we highlight the trade-off between fill probability itself and the subsequent returns of filled orders. We support our earlier assertions: first, that it is not at all easy to combine a high fill probability for maker orders with a positive subsequent return, due to one or other of adverse selection if filled, and low fill probability when subsequent returns are good; and second, for taker orders, to achieve short-term returns in excess of the taker fee. One can see this as an instance of an ‘Unprofitability Principle’ stated in cartoon form as
$\text{Ease of prediction}\times\text{ease of exploitation}<c$
for some positive constant $c$.
Adverse selection of maker orders has been studied in early works such as  foucault_2002 which discusses the need for market makers to monitor information sources to cancel stale quotes, while  foucault_2015 and  foucault_2016 examine the adverse selection risks market makers face from informed HFT takers reacting to news or arbitrageurs exploiting short-lived price adjustments. More recently,  DeLise:2024aa provided a more granular examination of adverse selection by analyzing post-fill price trajectories of maker orders. Our contribution extending these works lies in a precise description of the relationship between fill probability and post-fill returns at small timescales, and in demonstrating how this relationship shapes the way in which market makers should go about constructing their strategies.
@@ 1.1.3 Trading with Signals
We have noted (and analyze in more detail below) that some apparently natural strategies drawing on basic publicly available information (for example, posting a taker order in the direction of the imbalance or a maker order with high fill probability) may not yield a trading advantage. A trader may, however, take a more positive view if they have extra information in the form of a private signal, for example algorithmically generated.
Suppose, for example, that the top bid queue is large and the top ask queue is small, a situation we describe as a price-positive imbalance.  [^7] In markets with large relative tick sizes (in basis points) or substantial maker rebates such as ours, liquidity typically concentrates at the top bid and ask levels, and a high fraction of overall trading volume executes against them Menkveld2013. Intuitively, it is then likely that the shorter (ask) queue is the first to empty, in which case the spread widens and may be refilled with a new queue on the bid side, giving a price increase. This argument is corroborated in practice: it is well known that imbalance is positively correlated with the direction the next price movement. (For further insights, refer to  Gould:2015aa.) Now suppose that a trader supplements this modestly return-positive signal with a convincing private signal, suggesting an enhanced short-term return. How best to capitalize on this opportunity? Submitting a maker order at the end of the large bid queue is unlikely to result in a fill, and thus may not profit from the prediction (in the less frequent instances where it does fill, the short-term return is likely to be negative, as we discuss in more detail in Section 5). Alternatively, a taker order is certain to fill and benefit from any favorable price change; but they must pay the taker fee.
@@
These arguments illustrate the important point that the choice of order type and execution mode depends on the kind of signal in response to which the trader makes decisions. The literature contains numerous discussions of trading with signals, often under the umbrella term ‘optimal market making’ when they involve maker orders. For instance, Cartea2023 conclude that predictive signals, referred to as alpha signals, can boost market maker profitability. Their conclusions are arrived at under the assumption that maker orders fill upon the arrival of the next opposite-direction taker order; that is, the maker order is always at the front of the queue. As just argued, and discussed in more detail later, this is not generally tenable. Other papers such as  Fushimi2018OptimalHM and  Ait_Sahalia_2013 rely on the same assumption, while others still make different assumptions which are also problematic: for example, spooner_rl assumes that cancellations remove liquidity uniformly from the queue.
@@
Our examples above further illustrate the influence of queue position—whether submitting at the end of a long queue or a short one—on an order’s fill probability and the effectiveness of using maker orders to act on a trading signal. For example, posting maker orders in the direction of the imbalance implies posting at the back of the longer of the bid- and ask queues, unless one is the first to close the spread after a price change, which is extremely competitive and typically involves latency optimization and/or special order types. Conversely, one can obtain a good queue position (in the shorter of the two queues) by posting a maker order opposite to the imbalance-implied movement, but this entails trading against the expected return. Although queue position is rather sparsely addressed in the academic literature, Moallemi_2016 find that bad queue positions are associated with lower fill probabilities and increased adverse selection, findings that our results corroborate; additionally, Donnelly_2018 explore the interplay of queue position, adverse selection, and cancellation behavior, showing that within their modeling framework, the optimal strategy sometimes involves submitting an order counter to the imbalance to benefit from a queue position close to the front-of-queue, a result that is in agreement with our findings.
@@ 1.2  Notation and Definitions
Our analysis is set in a standard LOB, a data structure that records queues of maker orders waiting to be executed or canceled, prioritized by price and then by time of arrival, and quantized by the minimum order size, at price levels incremented by the tick size. The state of the order book at time $t$ is defined by the bid and ask sides:
@@
Its value is close to +1+1 when the liquidity is significantly greater at the top bid level than at the top ask level, and conversely for values near 1-1. As noted above, these extremes are associated with price increases and decreases respectively.
We focus exclusively on those maker orders posted at the top level (and taker orders filling them), meaning sell orders at the top ask price $p^{a}_{t,1}$ and buy orders at the top bid price $p^{b}_{t,1}$. These two order book levels hold a special significance as a large fraction of overall trading volume is executed against orders posted at the touch, making them especially important for market makers. For brevity, we shall use the term maker order to mean exclusively a maker order posted at one of the two top levels. The analysis of maker orders posted at deeper levels in the LOB requires different tools and is beyond the scope of this paper.
@@
Consider a buy order (and analogously for sell orders) $\mathfrak{o}$ posted at the top bid price $p^{b}_{t,1}$ at time $t$, indexed by $j_{0}$ in the sequence $\left(q^{b}_{t,1,j}\right)_{j}$ of individual order quantities. As illustrated in Figure 1, we define the liquidity ahead of the order as the total quantity of all preceding orders in the queue:
@@
and equal to 0 when there are no other orders in the queue ($\operatorname{LA}_{t}=\operatorname{LB}_{t}=0$). An order whose queue position is 0 is at the front of the queue; if it is $1$, the order is at the back.
@@
Consider an order $\mathfrak{o}$ (either maker or taker) posted at time $t_{0}$ when the top-of-book price is $p_{t_{0}}$.  [^8] Conventionally, for a maker order, $p_{t_{0}}$ is the top price on the near-side, while for a taker order, it is the top price on the opposite-side. We define a sign function, $\operatorname{sgn}(\mathfrak{o})$, equal to +1+1 for buy orders and 1-1 for sell orders, and multiply the drift values by this sign to obtain a direction-independent measure, with positive (negative) values indicating favorable (adverse) drift.